OpenStudy (anonymous):

I need help solving and writing a set builder notation for -3(x-10)>-x+5

OpenStudy (accessdenied):

You solve it by isolating the x by adding/subtracting the unrelated terms to the other side and then multiplying or dividing by the coefficient to get rid of it, obeying the rule that if you multiply/divide both sides by a negative, you switch the signs (its the only big difference between equality and inequality)

OpenStudy (anonymous):

I don't have to multiply -3 into (x-10) first?

OpenStudy (anonymous):

is proper to write set builder notion in fraction form or decimal?

OpenStudy (accessdenied):

you'd do that first, to make it easier in the long run...

OpenStudy (kinggeorge):

Usually I see set builder in fraction form, but it doesn't usually matter.

OpenStudy (anonymous):

okay thank you. Could I show how i've done the problem and you can let me know know I have the correct concept?

OpenStudy (accessdenied):

yeah, i think as long as you're consistent with which you choose, it's fine either way. fractions are usually nicer imo if you have repeating numbers though yeah, sure

OpenStudy (anonymous):

-30x-10>-x+5 30x>-x+5 30x+x>-x+x+5 30x>5 30x/5>5/5 ??

OpenStudy (accessdenied):

hmm... do you know how to solve a linear equation like 5x + 2 = 3x - 4? Where, you do the same thing to both sides to move the terms around?

OpenStudy (anonymous):

yes just learned

OpenStudy (accessdenied):

well, after we distribute the -3, we'd get... -3(x - 10) = -3(x) + (-3)(-10) = -3x + 30 -3x + 30 > -x + 5 We'd want to add and subtract to get our x-values on one side. +3x +3x 30 > 2x + 5 and then our constant terms on the other side -5 -5 25 > 2x then divide by the coefficient on x /2 /2 25/2 > x then rewrite this so that our x is first in the inequality x < 25/2

OpenStudy (anonymous):

Thank you so VERY much the set builder notation would be {x|x25/2} ?

OpenStudy (accessdenied):

you'd want the '<' in there between x and 25/2, but yeah, that would be correct.

OpenStudy (anonymous):

Really thank you so much that was my last question for class this week and I was stuck :)